Corecords has been growing at a rate of 34% per year in recent years. This same growth rate is expected to last for another 3 years, then to decline to gn = 6.0%.

If D1 = $0.6 and rs = 10.70%, what is the company's stock worth today?
Answer #1

HI

Here D1 = $0.6

growth rate for 3 years = 34%

so D2 = 0.6*(1+34%) = $0.80

D3 = 0.80*(1+34%) = $1.08

gn = 6%

rs = 10.70%

so stock worth today = D1/(1+rs) + D2/(1+rs)^2+D3/(1+rs)^3+D3*(1+gn)/(rs-gn)(1+rs)^3

=0.6/(1+10.70%) + 0.80/(1+10.70%)^2+ 1.08/(1+10.70%)^3+ 1.08*(1+6%)/(10.70%-6%)(1+10.70%)^3

= 19.90

Thanks